# Deep Reinforcement Learning for Bipedal Locomotion: A Brief Survey

### v1

@@ III-A Deep planning hybrid scheme
However, the limitations imposed by residual learning constrained the agents’ capacity to explore a broader array of possibilities. Building on previous work [25], a more efficient hybrid framework was developed, which learns from scratch without reliance on prior knowledge [43]. In this approach, a purely learning-based HL planner interacts with an LL controller using an Inverse Dynamics with Quadratic Programming formulation (ID-QP). This policy adeptly captures dynamic walking gaits through the use of reduced-order states and simplifies the learning trajectory. Demonstrating robustness and training efficiency, this framework has outperformed other models and was successfully generalized across various bipedal platforms, including Digit, Cassie, and RABBIT.
@@
In parallel, several research teams have focused on developing navigation planners specifically for toy-like humanoid robots, which provide greater physical stability compared to torque-driven or hydraulic bipedal robots as shown in Fig. 1. One notable study [46] implemented a visual navigation policy on the NAO robot, depicted in Fig. 1(a), utilizing RGB cameras as the primary sensory modality. This system has demonstrated successful zero-shot transfer to real-world scenarios, enabling the robot to adeptly navigate around obstacles. Further research [44] has explored complex dynamic motion tasks, such as playing soccer, by integrating a learned policy with an online footstep planner that utilises weight positioning generation (WPG) to create a center of mass (CoM) trajectory. This configuration is coupled with a whole-body controller, facilitating dynamic activities like soccer shooting. Despite their platform’s stability, provided by large feet and a lightweight structure, these robots exhibit limited dynamic movement capabilities compared to full-sized humanoid robots. Consequently, this research primarily addresses navigation and task execution.
@@
Regarding generalization, these frameworks have shown potential for adaptation across different types of bipedal and humanoid robots with minimal adjustments, demonstrating advanced user command tracking [43] and sophisticated navigation capabilities [44]. However, limitations are evident, notably the absence of capabilities for executing more complex and dynamic motions, such as jumping. Furthermore, while these systems adeptly navigate complex terrains with obstacles, footstep planning alone is insufficient without concurrent enhancements to the robot’s overall locomotion capabilities. Moreover, the requisite communication between the two distinct layers of the hierarchical framework may introduce system complexities. Enhancing both navigation and dynamic locomotion capabilities within the HL planner remains a significant challenge.
@@
Figure 3: Hierarchy Control Scheme Diagram: (a) A basic hierarchical scheme with two layers, where each module can be substituted with a learned policy. (b) A deep planning hybrid scheme, where the High-Level (HL) planner is learned. (c) A learning-based feedback control hybrid scheme, with a learned Low-Level (LL) controller. (d) A comprehensive DRL hierarchy control scheme, where both layers are learned.
@@ III-B Feedback DRL control hybrid scheme
For instance, gait libraries, which provide predefined movement references based on user commands, have been integrated into such frameworks [45]. Despite the structured approach of using gait libraries, their static nature offers limited adaptability to changing terrains, diminishing their effectiveness. A more dynamic approach involves online planning, which has shown greater adaptability and efficiency. One notable framework combines a conventional foot planner with an LL DRL policy [26], delivering targeted footsteps and directional guidance to the robot, thereby enabling responsive and varied walking commands. Moreover, HL controllers can provide additional feedback to LL policies, incorporating CoM or end-feet information, either from model-based methods or other conventional control strategies. However, this work has not yet been transferred from simulation to real-world applications. Later, a similar structure featuring an HL foot planner and an LL DRL policy was proposed [27]. This strategy not only achieved a successful sim-to-real transfer but also enabled the robot to navigate omnidirectionally and avoid obstacles.
@@
Compared to traditional model-based controllers, learned DRL policies provide a comprehensive closed-loop control strategy that does not rely on assumptions about terrain or robotic capabilities. These policies have demonstrated high efficiency in locomotion and accurate reference tracking. Despite their extensive capabilities, such policies generally require short inference times, making DRL a preferred approach in scenarios where robustness is paramount or computational resources on the robot are limited. Nonetheless, these learning algorithms often face challenges in environments characterized by sparse rewards, where suitable footholds like gaps or stepping stones are infrequent [48].
@@
Despite the significant potential demonstrated by previous studies, integrating DRL-based controllers with sophisticated and complex HL planners still presents limitations compared to more integrated frameworks such as end-to-end and deep planning models. Specifically, complex HL model-based planners often require substantial computational resources to resolve problems, rely heavily on model assumptions, necessitate extensive training periods, demand large datasets for optimization and hinder rapid deployment and iterative enhancements [48].
@@ III-C Learned hierarchy framework
The genesis of this framework was within a physics engine, aimed at validating its efficiency through simulation [12]. In this setup, LL policies, informed by human motions or trajectories generated via Trajectory Optimization (TO), strive to track these trajectories as dictated by the HL planner while maintaining balance. An HL policy is then introduced, pre-trained with long-term task goals, to navigate the environment and identify optimal paths. This structure enabled sophisticated interactions such as guiding a biped to dribble a soccer ball towards a goal. The framework was later enhanced to include imitation learning, facilitating the replication of dynamic human-like movements within the simulation environment [13].
However, despite its structured and layered approach, which allows for the reuse of learned behaviors to achieve long-term objectives, these frameworks have predominantly been validated only in simulations. The interface designed manually between the HL planner and the LL controller sometimes leads to suboptimal behaviors, including stability issues like falling.
@@
Within such a framework, the High-Level (HL) policy is optimized for strategic planning and achieving specific goals. This optimization allows for targeted enhancements depending on the tasks at hand. Moreover, the potential for continuous improvement and adaptation through further training ensures that the system can evolve over time, improving its efficiency and effectiveness in response to changing conditions or new objectives.
Despite the theoretical advantages, the practical implementation of this type of sim-to-real application for bipedal robots remains largely unexplored. The transition from simulation to real-world scenarios is fraught with challenges, not least because of the complexities involved in training and integrating two separate layers within the control hierarchy. Ensuring effective communication and cooperation between these layers is critical, requiring a meticulously defined communication interface to avoid operational discrepancies.
@@
Additionally, the training process for each policy within the hierarchy demands considerable computational resources. The intensive nature of this training can lead to a reliance on the simulation environment, potentially causing the system to overfit to specific scenarios and thereby fail to generalize to real-world conditions. This limitation highlights a significant hurdle that must be addressed to enhance the viability of learned hierarchy frameworks in practical applications.
IV Challenges and future research directions
While learning-based frameworks for bipedal robots have demonstrated considerable potential, they have also clearly exposed the limitations inherent to each framework. Moreover, several critical areas remain largely unexplored, especially within the realm of legged robotics, where the pace of research on bipedal robots lags behind that of their quadruped counterparts. This discrepancy in research progress can be attributed to several factors, including the higher costs and less mature technology associated with bipedal robot hardware, as well as the inherent instability issues that bipedal designs face.
To gain a deeper understanding of these challenges and to outline potential future directions, it is instructive to first review existing research on quadruped robots. The insights gained from quadrupeds, which benefit from more robust research outputs and technological advancements, can provide valuable lessons for overcoming similar challenges in bipedal systems.
IV-A Recent progress with quadruped robots
While DRL remains an emerging technology in bipedal robotics, it has firmly established its presence in the realm of quadruped robots, another category of legged systems. The diversity of frameworks developed for quadrupeds ranges from model-based RL designed for training in real-world scenarios, where unpredictable dynamics often prevail [49, 50], to systems that include the modeling of deformable terrain to enhance locomotion over compliant surfaces [51]. Furthermore, dynamic quadruped models facilitate highly adaptable policies [52], and sophisticated acrobatic motions are achieved through imitation learning [53].
The domain of quadruped DRL has also seen significant advancements in complex hybrid frameworks that integrate vision-based systems. To date, two primary versions of such frameworks have been developed: one where a deep planning module is paired with model-based control [54], and another that combines model-based planning with low-level DRL control [48, 55]. The latter has shown substantial efficacy; it employs a model predictive control (MPC) to generate reference motions, which are then followed by a LL feedback DRL policy. Additionally, the Terrain-aware Motion Generation for Legged Robots (TAMOLS) module [56] enhances the MPC and DRL policy by providing terrain height maps for effective foothold placements across diverse environments, including those not encountered during training. However, similar hybrid control schemes have not been thoroughly investigated within the field of bipedal locomotion.
Despite the progress in quadruped DRL, similar advancements have been limited for bipedal robots, particularly in loco-manipulation tasks and vision-based DRL frameworks. Establishing a unified framework could bridge this gap, an essential step given the integral role of bipedal robots in developing full humanoid systems. Moreover, the potential of hybrid frameworks that combine model-based and DRL-based methods in bipedal robots remains largely untapped.
IV-B Gaps and challenges
Despite numerous promising developments in the field of bipedal and humanoid robotics, significant gaps remain between current research outcomes and the ultimate goals. This discussion concentrates on the gaps in frameworks and algorithms rather than hardware, structured around two pivotal questions: 1) Is it possible to design a unified framework that achieves both generalization and precision? 2) Can we develop a straightforward end-to-end policy capable of managing all tasks efficiently?
IV-B1 Generalization versus precision
@@
These challenges underscore the need for innovative solutions that can bridge the gap between the capabilities of current frameworks and the ambitious goals of advanced bipedal and humanoid robotics.
IV-B2 Simplifying frameworks to overcome complex tasks
The envisioned ideal in robotic design is an end-to-end framework that enables robots to traverse various terrains using versatile locomotion skills. Although current research often focuses on enhancing frameworks by adding complex components to mitigate inherent limitations, such as the integration of a foot planner for omnidirectional locomotion and stair navigation, as demonstrated in [26, 27], simpler end-to-end frameworks have also proven effective. These frameworks adeptly navigate challenging terrains and perform a diverse range of locomotion tasks with fewer components [20, 22].
The advantage of maintaining simplicity in the framework lies in its ability to streamline decision-making processes, thereby reducing computational overhead and potential points of failure. To achieve an optimal end-to-end framework, advancements in several key areas are essential:

### v3

@@ III-A Deep planning hybrid scheme
However, the limitations imposed by residual learning constrained the agents’ capacity to explore a broader array of possibilities. Building on previous work [34], a more efficient hybrid framework was developed, which learns from scratch without reliance on prior knowledge [53]. In this approach, a purely learning-based HL planner interacts with an LL controller using an Inverse Dynamics with Quadratic Programming formulation (ID-QP). This policy adeptly captures dynamic walking gaits through the use of reduced-order states and simplifies the learning trajectory. Demonstrating robustness and training efficiency, this framework has outperformed other models and was successfully generalised across various bipedal platforms, including Digit, Cassie, and RABBIT.
@@
In parallel, several research teams have focused on developing navigation planners specifically for toy-like humanoid robots, which provide greater physical stability compared to torque-driven or hydraulic bipedal robots as shown in Fig. 1. One notable study [57] implemented a visual navigation policy on the NAO robot, depicted in Fig. 1(a), utilising RGB cameras as the primary sensory modality. This system has demonstrated successful zero-shot transfer to real-world scenarios, enabling the robot to adeptly navigate around obstacles. Further research [54] has explored complex dynamic motion tasks, such as playing soccer, by integrating a learned policy with an online footstep planner that utilises weight positioning generation (WPG) to create a centre of mass (CoM) trajectory. This configuration is coupled with a whole-body controller, facilitating dynamic activities like soccer shooting. Despite their platform’s stability, provided by large feet and a lightweight structure, these robots exhibit limited dynamic movement capabilities compared to full-sized humanoid robots. Consequently, this research primarily addresses navigation and task execution.
@@
Regarding generalisation, these frameworks have shown potential for adaptation across different types of bipedal robots with minimal adjustments, demonstrating advanced user command tracking [53] and sophisticated navigation capabilities [54]. However, limitations are evident, notably the absence of capabilities for executing more complex and dynamic motions, such as jumping. Furthermore, while these systems adeptly navigate complex terrains with obstacles, footstep planning alone is insufficient without concurrent enhancements to the robot’s overall locomotion capabilities. Moreover, the requisite communication between the two distinct layers of the hierarchical framework may introduce system complexities. Enhancing both navigation and dynamic locomotion capabilities within the HL planner remains a significant challenge.
@@
Figure 3: Hierarchy Control Scheme Diagram: presents a hierarchical control framework for a bipedal robot, consisting of a basic scheme and three variations. (1) Basic Scheme: The framework starts with a Task Command, followed by a HL Planner, and a LL Controller, which ultimately drives the Robot. Each module can be replaced by a learned policy, introducing adaptability in different control layers. (2) Variations (from left to right): (a) A deep planning hybrid scheme, where the HL planner is learned. (b) A Feedback DRL control hybrid scheme, with a learned LL controller. (c) A Learned hierarchy control scheme, where both layers are learned.
@@ III-B Feedback DRL control hybrid scheme
For instance, gait libraries, which provide predefined movement references based on user commands, have been integrated into such frameworks [55]. Despite the structured approach of using gait libraries, their static nature offers limited adaptability to changing terrains, diminishing their effectiveness. A more dynamic approach involves online planning, which has shown greater adaptability and efficiency. One notable framework combines a conventional foot planner with an LL DRL policy [35], delivering targeted footsteps and directional guidance to the robot, thereby enabling responsive and varied walking commands. Moreover, HL controllers can provide additional feedback to LL policies, incorporating CoM or end-feet information, either from model-based methods or other conventional control strategies. However, this work has not yet been transferred from simulation to real-world applications. Later, a similar structure featuring an HL foot planner and an LL DRL-based policy was proposed [36]. This strategy not only achieved a successful sim-to-real transfer but also enabled the robot to navigate omnidirectionally and avoid obstacles.
@@
Compared to traditional model-based controllers, learned DRL policies provide a comprehensive closed-loop control strategy that does not rely on assumptions about terrain or robotic capabilities. These policies have demonstrated high efficiency in locomotion and accurate reference tracking [59]. Despite their extensive capabilities, such policies generally require short inference times, making DRL a preferred approach in scenarios where robustness is paramount or computational resources on the robot are limited. Nonetheless, these learning algorithms often face challenges in environments characterised by sparse rewards, where suitable footholds like gaps or stepping stones are infrequent [59].
@@
Despite the significant potential demonstrated by previous studies, integrating DRL-based controllers with sophisticated and complex HL planners still presents limitations compared to more integrated frameworks such as end-to-end and deep planning models. Specifically, complex HL model-based planners often require substantial computational resources to resolve problems, rely heavily on model assumptions, necessitate extensive training periods, demand large datasets for optimisation and hinder rapid deployment and iterative enhancements [59].
@@ III-C Learned hierarchy framework
The genesis of this framework was within a physics engine, aimed at validating its efficiency through simulation [21]. In this setup, LL policies, informed by human motions or trajectories generated via TO, strive to track these trajectories as dictated by the HL planner while maintaining balance. An HL policy is then introduced, pre-trained with long-term task goals, to navigate the environment and identify optimal paths. This structure enabled sophisticated interactions such as guiding a biped to dribble a soccer ball towards a goal. The framework was later enhanced to include imitation learning, facilitating the replication of dynamic human-like movements within the simulation environment [22].
However, despite its structured and layered approach, which allows for the reuse of learned behaviours to achieve long-term objectives, these frameworks have predominantly been validated only in simulations. The interface designed manually between the HL planner and the LL controller sometimes leads to suboptimal behaviours, including stability issues like falling.
@@
To fully leverage its potential, HumanPlus has been developed as a versatile framework for humanoid robots, integrating hierarchical learning, multimodal perception, and real-world imitation [56]. It employs a two-layer structure, where HIT learns from human demonstrations, trained on AMASS, and HST acts as a LL tracking controller. Additionally, binocular RGB vision input enhances perception, enabling precise loco-manipulation and dynamic locomotion tasks such as jumping, walking, folding clothes, and rearranging objects. This shadowing-based imitation learning approach improves adaptability, making it a promising framework for transferring human-like skills to robots.
@@
Within such a framework, HL policy is optimised for strategic planning and achieving specific goals. This optimisation allows for targeted enhancements depending on the tasks at hand. Moreover, the potential for continuous improvement and adaptation through further training ensures that the system can evolve over time, improving its efficiency and effectiveness in response to changing conditions or new objectives. Despite the theoretical advantages, the practical implementation of this type of sim-to-real application for bipedal robots remains largely unexplored.
@@
Additionally, the training process for each policy within the hierarchy demands considerable computational resources [38]. The intensive nature of this training can lead to a reliance on the simulation environment, potentially causing the system to overfit to specific scenarios and thereby fail to generalise to real-world conditions. This limitation highlights a significant hurdle that must be addressed to enhance the viability of learned hierarchy frameworks in practical applications.
Besides, for the general hierarchical framework, the transition from simulation to real-world scenarios is challenging, particularly due to the complexities involved in coordinating two layers within the control hierarchy. Ensuring seamless communication and cooperation between the HL Planner and LL Controller is essential to avoid operational discrepancies. The primary challenges include: (1) Task division complexity—while the HL Planner handles strategy and provides abstract goals, the LL Controller manages precise execution, necessitating careful coordination to avoid functional overlap and conflicts. (2) Effective communication—the HL’s abstract goals must be accurately interpreted and converted by the LL into real-time actions, especially in dynamic environments. (3) Task allocation—clear division of responsibilities between layers is crucial to prevent redundancy and ensure smooth system performance.
IV Towards a Unified Framework
Previous Section II and Section III illustrate two distinct approaches for controlling bipedal robots. Whether it is end-to-end or hierarchical, their ultimate goal is to adeptly navigate all types of terrains Both these two finally towards to the unified frameworks, can be named as Humanoid Foundation Model (HFM) and Multi-layered Adaptive Models (MLAM), which are expanded based on end-to-end, and hierarchical approaches.
In this section, we explore the conceptual design of HFM and MLAM, analyse the existing gaps and challenges in achieving these unified frameworks, and discuss recent advancements in quadruped robotics that offer valuable insights for the future development of bipedal locomotion.
IV-A Humanoid Foundation Models
Inspired by the Robot Foundation Models (RFMs) [60, 61], built on an end-to-end framework, enhance versatility by learning shared representations across different skills, HFM focus specifically on humanoid robots, which inherit advantages of RFMs but possess a higher number of DoFs and a human-like morphology, requiring more advanced motion control and adaptability. As foundation model, this structure bypasses the need for task decomposition, allowing the system to learn through large datasets and generalise across diverse locomotion tasks.
HFM with large parameters are expected to be pre-trained on large-scale robotic datasets across various platforms and later fine-tuned using DRL to adapt to specific locomotive tasks. This enables humanoid robots to continuously refine their locomotion strategies and adapt to dynamic environments through interaction. For instance, AdA [62] demonstrates a framework where an agent is first pre-trained on extensive datasets and subsequently fine-tuned with RL, allowing it to optimise policy performance and achieve near-optimal control.
Other foundation model work also provide thoughts about how to build up HFM. For example Current research on RFMs mainly about Vision-Language-Action (VLA) model on robotic arm platform, derived from VLM, VLA models directly map visual information to robot actions, enabling end-to-end decision-making without explicit intermediate representations. Recent work, such as RT-2 [63], expands upon the transformer model-based RT-1 [64], a vision-language model designed for robotic control. In RT-2, the entire model is trained using a combination of a VLM and robotic demonstrations. The actions need to be quantized as inputs and subsequently dequantized.
Challenges: The challenges of RMFs can be derived from two aspects: (1) training datasets and (2) model inference time.
  1. 1.
    Training datasets: In terms of datasets, unlike Large Language Models (LLMs) or VLMs, HMF takes multimodal inputs with heterogeneous data formats that are not inherently aligned. While LLMs process textual prompts as both input and output, and VLMs typically map images and text bidirectionally, HFM for bipedal robot must handle diverse sensory data, including vision, proprioception, motion trajectories, and object interactions, making input representation and model alignment more complex. Which lead to input representation and feature alignment challenge, which means different data modalities have different feature distributions and dimensionalities, making direct fusion difficult.
    Beside of directly making the training process difficult and high requirements of model construction, another challenge in datasets is datasets collection. Current datasets provide human motions for bipedal humanoid robot to imitate. For example, AMASS [65] consists solely of human motion capture data, while Motion-X [66] extends this by incorporating textual prompt labels corresponding to the motions. Although various references exist for robots to learn different skills, most datasets are not directly tailored for robots and require additional mapping to specific robotic platforms.
  2. 2.
    Model inference time: Bipedal robots require high-frequency action strategies to maintain stability, but real-time decision-making in HFM faces several limitations. First, RMFs using LLMs as backbones are computationally heavy, leading to slow inference due to expensive forward propagation. Second, autoregressive token generation introduces sequential dependencies, preventing parallel computation and further increasing latency. Third, onboard processors lack sufficient memory and computational power to run large models efficiently, making real-time execution challenging without model compression or hardware acceleration.
IV-B Multi-layered Adaptive Model
Unlike HFM, the MLAM follows a hierarchical framework, decomposing tasks into multiple sub-tasks. Each layer is responsible for a specific level of decision-making, with higher layers handling HL planning and lower layers executing fine-grained control. This structured approach enables the system to adapt dynamically and ultimately achieve the overall goal. Each module in the hierarchical model can be fine-tuned, replaced, or upgraded without requiring re-training of the entire model.
The discrete model is structured into three hierarchical levels: strategy, planning, and control, each progressively refining task execution. Section III has demonstrated how DRL-based policies contribute to both the HL planner and LL controller. At the highest level, the strategy layer serves as the reasoning component, enabling the robot to comprehend tasks, infer objectives, and decompose long-term goal-based tasks into structured sub-tasks, which are then passed down for execution. For example, [67] employs an LLM as a reasoning layer to interpret user commands, break them into sequential task steps, and translate them into executable code for a quadruped robot.
Challenge: In the previous Section LABEL:Sec:learned_heirarchical_framework, the challenges of the hierarchical framework were briefly discussed. However, when considering the unified framework—MLAM, additional challenges emerge that require further exploration. These challenges are:
  1. 1.
    Limited Visual Perception and Multi-Modal Integration: MLAM relies on real-time perception and decision-making, yet bipedal locomotion demands low-latency, high-precision sensory processing (e.g., vision, IMU, force sensors). Integrating these multi-modal inputs across different layers of the hierarchy remains a significant challenge. Each modality operates at different timescales and levels of abstraction, making seamless coordination and alignment difficult. Moreover, existing methods often fail to effectively capture cross-modal dependencies, limiting their applicability in dynamic scenarios.
  2. 2.
    Computational Efficiency and Real-Time Execution: MLAM must balance the quality of HL decision-making with the strict low-latency requirements of real-time control. This is especially problematic for HL reasoning models, such as LLMs, which are inherently designed for step-by-step reasoning and rely on autoregressive decoding. Generating decisions token-by-token significantly increases inference time, making these models unsuitable for tasks that demand rapid response times. These delays can lead to unstable behaviours, particularly in tasks requiring fast adaptation and precise coordination in dynamic environments.
These challenges highlight the complexity of implementing MLAM and underscore the need for innovative solutions to address these limitations in real-world applications.
IV-C Fundamental challenges
The development of unified frameworks such as HFM and MLAM aims to address the challenges of bipedal locomotion in dynamic and unstructured environments. However, achieving such frameworks requires overcoming several fundamental obstacles. Among these, balancing generalisation and precision, as well as designing streamlined frameworks capable of handling complex tasks, represent two critical challenges. These challenges are explored in the following subsections.
IV-C1 Generalisation versus precision
The challenge of balancing generalisation and precision is a fundamental obstacle in achieving the unified frameworks of HFM and MLAM. For HFM, which focuses on large-scale, versatile humanoid models, the difficulty lies in enabling the model to adapt to diverse tasks while maintaining precise control in dynamic environments. For MLAM, which emphasizes task decomposition and hierarchical control, ensuring precision across sub-tasks while maintaining a HL general policy is equally challenging.
@@ IV-C1 Generalisation versus precision
For HFM, these challenges are exacerbated by its reliance on diverse and large-scale datasets, which often lack the precision required for specific bipedal tasks. In contrast, MLAM struggles to achieve precision across its hierarchical layers, where HL policies must generalise across tasks while LL controllers maintain stability and accuracy in dynamic scenarios.
@@
These challenges underscore the need for innovative solutions that balance task generalisation with control precision, bridging the gap between HFM’s large-scale adaptability and MLAM’s hierarchical precision.
IV-C2 Streamlined frameworks capable of handling complex tasks
One of the critical challenges in unifying HFM and MLAM is designing frameworks that are both simple and capable of handling complex tasks. While streamlined frameworks reduce computational overhead and potential failure points, they may lack the flexibility to adapt to diverse and dynamic environments. On the other hand, complex frameworks with multiple components, such as foot planners or terrain-specific modules [35, 36], often introduce additional computational burden and coordination challenges.
For HFM, the challenge lies in creating a model that leverages large-scale general representations without adding unnecessary layers of complexity. For MLAM, the difficulty is in coordinating hierarchical decision-making across layers without introducing redundancy or inefficiency. Achieving a balance between simplicity and capability is essential for creating robust, unified frameworks.
The advantage of maintaining simplicity in the framework lies in its ability to streamline decision-making processes, thereby reducing computational overhead and potential points of failure. To achieve an optimal framework, advancements in several key areas are essential: